OpenStudy (anonymous):

can anyone teach me how to graph these functions? f(x+3)+2 f(x-3)-2 f(x+2)+3 f(x-2)-3

OpenStudy (campbell_st):

so f(x + 3) is function notation you need a function f(x) = ...... do you have it..?

OpenStudy (zale101):

As Campbell states, we can't graph it because we don't know what f(x) is. It's just stating the basic shifting of whatever the f(x) graph is. f(x-h)+k where h is the x value and k is the y value +K shifts up, if it was negative (-K), it'll shift down f(x-h) shifts to the right because h is f(x-(+h)), it will shift to the left if h is negative f(x-(-h)) which is f(x+h)

OpenStudy (anonymous):

@zale101 @campbell_st the parent function is f(x)=log2 x

OpenStudy (campbell_st):

well a simple solution is to use some graphing software plot \[y = \log_{2} (x)\] then on the same set of axes plot \[y = \log_{2} (x + 3) + 2\] look at the effect of adding 3 to x does... then study the effect of + 2 use this site https://www.desmos.com/calculator

OpenStudy (anonymous):

@campbell_st thanks a lot, can you please help me with one more question?

OpenStudy (campbell_st):

sure, post it

OpenStudy (anonymous):

which logarithmic graph can be used to approximate the value of y in the equation 6^y=12?

OpenStudy (anonymous):

the first graph again

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

well if you take the log of both sides its \[\ln(6^y) = \log(12)\]apply the log law for powers \[ylog(6) = \log(12)\] so \[y = \frac{\log(12)}{\log(6)}\] so looking at this y is a positive value larger than 1. there appears only 1 graph that has a positive y value

OpenStudy (anonymous):

but all of the graphs (from what i see) have negative y values :(

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

which one has a point above the horizontal axis...

OpenStudy (anonymous):

the first and second graph both have points above the horizontal axis

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

ok... so do the calculation ln(12)/ln(6) = 1.386.. so which has a point at (12, 1.386)

OpenStudy (anonymous):

is it the first graph?

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

I think the 1st is the better choice... the 2nd looks like about 1

OpenStudy (anonymous):

alright thanks can you help me with one last question. I am having doubts with it really, i want to see if i'm right @campbell_st

OpenStudy (anonymous):

The function f(x)=245(4)^x represents the growth of a fruit fly population every YEAR in the orange grove. Stacy wants to manipulate the formula to an equivalent form that calculates every MONTH, not every year. Which function is correct for Stacy's purposes? f(x)=45(4)^x f(x)=45(4^12)x/12 f(x)=2752(4)^x f(x)=245(4^1/12)^12x II think that it is the second option but idk. My reasoning is because since in a single year there are 12 months then it should be the denominator of a fraction and the numerator would be howm many months. But im also not sure if it might be the last one

OpenStudy (campbell_st):

sorry, just post it in the normal way, someone will help